DOPPLER-FREE STUDY OF THE TiO MOLECULE

C. AMIOT, P. LUC, and \underline{R. VETTER}, Laboratoire Aimé Cotton (CNRS UPR3321), Campus d’Orsay, Bâtiment 505, 91405 Orsay, France.

We have studied the yellow-orange absorption spectrum of the TiO molecule by use of a Doppler-free spectroscopy experiment in which a cw tunable laser beam crosses an effusive beam of titanium monoxide. Thousands of wavenumbers belonging to the B^{3}Π - X^{3}Δ(1-0) and c^{1}Φ - a^{1}Δ(0-0) bands have been measured to an accuracy of 10^{−3} cm^{−1} and interpreted in terms of molecular constants. For the first band^{a}, analysis of fine structure and Λ-doubling shows that the electronic structure of the B^{3}Π state must be revisited. For the second band^{b}, electronic interactions with distant vibrational levels of the c^{3}Δ state must be considered to assign high rotational levels. Analysis of the spectra due to the five isotopes of titanium is on the tracks.

aC. Amiot, E. M. Azaroual, P. Luc and R. Vetter, J. Chem.Phys., 102, 4375 (1995).

bC. Amiot, M. Cheikh, P. Luc and R. Vetter, J. Mol. Spectroscopy , 179,159 (1996).